### v1

# Critical states exhibit invariance in both position and momentum spaces

Abstract
The critical states of disordered systems are intriguing subjects within the realm of condensed matter physics and complex systems. These states manifest in materials where disorder plays a significant role, and are distinguished by their multifractal structure and self-similarity. However, accurately characterizing critical states continues to pose a significant challenge. In this study, we argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains. More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space. Subsequent numerical simulations validate the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
I Introduction
The concept of critical states Critical in various physical systems pertains to the behaviors and characteristics exhibited during phase transitions or as they approach critical points. The examination of critical states provides profound insights into the nature of order, disorder, and collective behavior across diverse disciplines Sachdev1 . These states are fundamental for comprehending a wide range of phenomena in statistical physics Critical0 , condensed matter physics Critical1 , and complex systems Critical2 . As a defining feature of critical phenomena Critical3 , fluctuations in critical states become correlated over large distances, with the system demonstrating self-similar behavior at different scales. Critical fluctuations have been observed in various systems, including phase transitions in ferromagnets-antiferromagnets Critical4 , percolation thresholds in complex networks Critical5 , and photonic quasicrystals photonic . As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of complex systems.
Recently, significant advancements have been achieved in the investigation of critical states of wave functions, specifically focusing on critical states of quantum disordered systems Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; gg4 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 . Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging. Nevertheless, the primary challenge in comprehending critical states lies in determining the conditions under which they arise and accurately characterizing them.
Bohr’s complementarity principle MQM provides us with inspiration to bridge this gap. According to the complementarity principle, it is impossible to simultaneously observe the wave nature and particle nature of matter under different representations, such as position space and momentum space. Matter exhibits different properties in these representations. The relationship between position space and momentum space is primarily established through the Fourier transform; position space serves as a mathematical framework for describing the location of particles or objects in physical space, while momentum space is perceived as the duality of position space QM . The critical state can be considered as a third state in addition to particle and wave properties, thus it can be reasonably inferred that the critical state may exhibit some invariance in both position space and momentum space.
Figure 1: (Color online) (a) and (b) demonstrate an application of Bohr’s complementarity principle, showing that a localized state in position space must be delocalized in momentum space, and vice versa for an extended state. (c) and (d) illustrate that critical states can simultaneously exhibit delocalization, multifractality, and self-similarity in both position space and momentum space.
Based on this duality perception, Liu and Xia propose an explicit criterion Liu-Critical for precisely characterizing critical states. They assert that the Lyapunov exponents of critical states should simultaneously be 0 in both position space and momentum space. The criterion definitively distinguishes critical states from extended and localized states, as the Lyapunov exponents of the latter two are not equal in position space and momentum space. If one exponent is 0, then the dual exponent must be greater than 0. In physics, this also implies that extended and localized states cannot simultaneously exhibit delocalization in position space and momentum space, as illustrated in Fig. 1(a) and (b). In contrast, critical states should demonstrate delocalization, multifractality, and self-similarity in both position space and momentum space, as depicted in Fig. 1(c) and (d). These findings clearly illustrate that critical states maintain a certain kind of invariance across two dual spaces.
A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces? This study demonstrates that the invariance of the critical state is universal, encompassing various physical quantities such as the inverse participation ratio and information entropy.
II Analytical derivation
The application of Lyapunov exponents to disordered systems Xia ; Jitomirskaya ; Avila ; Dwiputra focuses on calculating the decay rate of the wave function. For a given disordered potential, the conjectured solution for the wave function in the $n$-th lattice is $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$, where $\gamma\geq 0$ represents the Lyapunov exponent. A value of $\gamma=0$ indicates that $|\psi_{n}|\sim|\psi_{0}$, corresponding to an extended state; while a value of $\gamma>0$ indicates that $|\psi_{n}|\sim|\psi_{0}|e^{-\gamma n}$, corresponding to a localized state. This proposed solution provides a reasonable framework encompassing eigenstates of disordered systems and can be utilized to derive mathematical expressions for the Inverse Participation Ratio and information entropy with respect to $\gamma$.
The Inverse Participation Ratio (IPR) is a physical quantity IPR ; IPR1 ; IPR2 commonly utilized in the fields of quantum mechanics and condensed matter physics to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 , particularly within disordered systems. It serves as a quantitative measure of the extent to which a wave function is distributed across space versus its degree of localization. A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
Information entropy $S$, commonly known as Shannon entropy Shannon in honor of Claude Shannon, is a metric for the unpredictability associated with a system. Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum complex systems IE1 . Information entropy has been utilized to comprehend and quantify the level of unpredictability and randomness in disordered systems IE2 ; IE3 .
The physical quantities IPR and $S$ can be explicitly expressed using the wave function Ansatz $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$. By applying the normalization condition $\sum_{n=1}^{L}|\psi_{n}|^{2}=1$, we can deduce the form of the initial wave function $|\psi_{0}|^{2}\sim e^{2\gamma}-1$. Then utilizing the definition of IPR, we can get
${\rm IPR}=\lim_{n\rightarrow\infty}\sum_{n=1}^{L}|\psi_{n}|^{4}\sim\frac{e^{2\gamma}-1}{e^{4\gamma}-1}\sim\tanh(\gamma).$ (1)
In the same manner, we can establish its relation with Lyapunov exponent by utilizing the definition of information entropy,
$S=-\lim_{n\rightarrow\infty}\sum_{n=1}^{L}|\psi_{n}|^{2}\ln(|\psi_{n}|^{2})\sim\gamma+\gamma\coth(\gamma)-\ln(-1+e^{2\gamma}).$ (2)
According to the Liu-Xia criterion Liu-Critical , critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. This can be mathematically expressed as the equality of the Lyapunov exponent of position space $\gamma$ and that of momentum space $\gamma_{m}$,
$\gamma=\gamma_{m}.$ (3)
From Eq. (1) and Eq. (2), it is evident that the IPR and information entropy $S$ both exhibit monotonic behavior with respect to the Lyapunov exponent. Therefore, according to Eq. (3), the IPR and $S$ of the critical state should be equal in both position space and momentum space, namely
${\rm IPR}={\rm IPR_{m}},~{}~{}~{}~{}S=S_{m}.$ (4)
Equation (4) represents the key innovation of this study, demonstrating the universal invariance of the critical state in both position space and momentum space. This concept applies not only to Lyapunov exponent, but also to other fundamental physical quantities such as IPR and $S$.
III Numerical verification
To validate the theoretical findings, we conduct numerical simulations to confirm the critical states in two typical quasiperiodic models. These models are chosen due to their easily expressible Hamiltonian in both position and momentum spaces.
III.1 Aubry-André-Harper model
The Aubry-André-Harper model AAH1 ; AAH2 (AAH) model describes a tight-binding system of non-interacting particles on a one-dimensional lattice, providing a description of electron behavior in quasiperiodic potentials. Introduced by S. Aubry and G. André in 1980, it has become a cornerstone in the study of quasiperiodic systems and localization phenomena. The Hamiltonian in position space can be expressed as a discrete Schrödinger equation:
$\psi_{n+1}+\psi_{n-1}+V\cos(2\pi\alpha n+\theta)\psi_{n}=E\psi_{n}.$ (5)
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential at site $n$, and $E$ is the energy eigenvalue of the particle. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity. $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity). Thus, the potential $V_{n}$ at each lattice site $n$ is quasiperiodic because it is a non-repeating function with long-range correlations, unlike random disordered potentials that are purely random at each site.
The AAH model yields an intriguing result in the form of a phase transition between two distinct phases as the strength of the quasiperiodic potential $V$ is varied. When $V$ is small, the wave functions extend over the entire system, leading to metallic behavior as particles can delocalize and move freely. In contrast, for large $V$, the system enters the Anderson localized phase where wave functions are spatially localized due to the strong potential, restricting particle propagation.
The phase transition occurs at $V=2$, where all eigenstates at this point are critical and display self-similarity. Utilizing the Fourier transform $\psi_{n}=\sum_{k}e^{-i2\pi\alpha nk}\phi_{k}$, the Hamiltonian of the AAH model in momentum space can be readily expressed,
$\frac{V}{2}(\phi_{k+1}+\phi_{k-1})+2\cos(2\pi\alpha k+\vartheta)\phi_{k}=E\phi_{k}.$ (6)
When $V=2$, Eq. (5) and Eq. (6) exhibit identical forms, thus $V=2$ is referred to as the self-dual point. The concept of self-duality holds significance due to its preservation of invariance under a Fourier transform, which interchangeably swaps the roles of position and momentum within the Hamiltonian.
III.2 Quasiperiodic-Nonlinear-Eigenproblem model
Liu and Xia propose a pioneering model Liu-Critical , which is the first nonlinear eigenvalue problem in Non-Hermitian physics. Let us explain what a linear or nonlinear eigenvalue problem is, the mathematical form of which can be expressed as
$\hat{H}\left|\psi\right\rangle=E\hat{B}\left|\psi\right\rangle,$ (7)
where $\hat{H}$ represents the Hamiltonian and $E$ is the eigenvalue. When the matrix $\hat{B}$ equals the identity matrix $I$, Eq. (7) corresponds to a linear eigenvalue problem, such as the well-known Schrödinger equation. On the other hand, when $\hat{B}\neq I$, Eq. (7) corresponds to a nonlinear eigenvalue problem or a generalized eigenvalue problem. Based on this concept, the Quasiperiodic-Nonlinear-Eigenproblem (QNE) model proposed by Liu and Xia Liu-Critical is formulated as
$\{2\cos[2\pi\alpha(n+1)]+V\}{\psi_{n+1}}+$ (8)
$\{2\cos[2\pi\alpha(n-1)]-V\}\psi_{n-1}=E(\psi_{n+1}+\psi_{n-1}),$
where the meaning of parameters is the same as that of the AAH model. The QNE model [Eq. (8)] differs significantly from the AAH model [Eq. (5)], particularly on the right-hand side of the equation, making it no longer a linear eigenvalue problem like previous quasiperiodic systems. The QNE model holds a unique position in the field of Non-Hermitian physics NH1 ; NH2 ; NH3 ; NH4 ; NH5 ; NH6 ; NH7 , because non Hermitian can be introduced in three ways: through non-reciprocity in hopping terms, through complex momentum involving gain and loss of reciprocal hopping terms, or by introducing complex mass representing gain and loss of on-site potential. However, the QNE model does not fall into any of these three cases as it is not a linear eigenvalue problem.
Therefore, the QNE model exhibits at least two non-trivial properties: (i) through the introduction of nonlinear eigenvalue terms, the system becomes non-Hermitian, distinguishing it from previous non-Hermitization methods such as complexification of hopping terms and on-site potential, or non-reciprocity of neighbor hoppings; (ii) critical states in the QNE model exist within a wide range of parameters $0<V\leq 2$, unlike the AAH model where critical states only emerge at the phase transition point $V=2$. Upon applying Fourier transform, the Hamiltonian of the QNE model in momentum space can be expressed as
$\phi_{k+1}+\phi_{k-1}+iV\tan(2\pi\alpha k)\phi_{k}=E\phi_{k}.$ (9)
where $i$ represents the imaginary unit, and disregarding the phase factor. Fortunately, diagonalizing the Hamiltonian matrix of the QNE model in momentum space results in a linear eigenvalue problem, significantly simplifying the numerical simulation process.
Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of critical states, IPR and $S$ in position space and momentum space for the AAH model, respectively. (d), (e) and (f) demonstrate the comparison of critical states, IPR and $S$ in position space and momentum space for the QNE model, respectively. Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. In the AAH model, IPR and $S$ are only equal between position space and momentum space at the phase transition point $V=2$, whereas in the QNE model, these physical quantities are equal within a large parameter range of $0<V\leq 2$.
III.3 Verification results
We perform direct numerical diagonalization on the AAH model and QNE model, obtaining the eigenvalues and eigenstates separately. The total number of lattices in the system is set to 1000. For the AAH model, our focus is on the middle eigenstate; whereas for the QNE model, we concentrate on the eigenstate with an eigenvalue of 0.
The numerical results are summarized in Fig. 2. For both the AAH model and the QNE model, the critical state in position space $|\psi|$ and momentum space $|\phi|$ exhibit delocalized, multifractal, and self-similar characteristics simultaneously, as depicted in Fig. 2 (a) and (d). In the AAH model, equality between physical quantities of position space and momentum space only occurs at the phase transition point $V=2$, specifically $\text{IPR}=\text{IPR}_{m}$ and $S=S_{m}$, as shown in Fig. 2 (b) and (c). However, for the QNE model, this equality is observed over a wider parameter range when $0<V\leq 2$, as illustrated in Fig. 2 (e) and (f).
The validation of previous theoretical predictions through numerical results is crucial in demonstrating the reliability of the invariance of critical states. These results confirm that essential physical quantities, such as the IPR and information entropy, exhibit invariance in both position space and momentum space at critical states. Moreover, the persistence of the previous Lyapunov exponent’s invariance further emphasizes the broad applicability of critical state invariance to a wide range of physical quantities.
This universal invariance across different physical quantities suggests that the stability of critical states transcends specific system details or characteristics. Whether dealing with low-dimensional or high-dimensional systems, Hermitian or non-Hermitian matrices, random disorder or non-random disorder, the concept of critical state invariance remains consistent. This observation is consistent with the perspective of Liu-Xia criterion Liu-Critical , which emphasizes that the invariance of critical states does not depend on specific system attributes but rather reflects a broader principle applicable across diverse scenarios. Furthermore, this concept may also be associated with conformal invariance, suggesting potential connections between the stability of critical states and fundamental principles governing symmetry and transformation properties within physical systems.
IV Summary
In conclusion, we have addressed a long-standing unresolved issue regarding the definition of a critical state and its accurate characterization. Our findings emphasize the universality of critical state invariance across various domains in physics.
The critical states in quantum disordered systems have been a subject of extensive research due to their unique multifractal and self-similar structure of wave functions. Despite the significant interest in these states, experimental manipulation has proven to be a highly challenging task. In this study, we aim to provide a comprehensive characterization of the universal properties of critical states. One key finding is that physical quantities associated with critical states remain invariant in both position space and momentum space. This discovery fills a theoretical gap in our understanding of how to precisely characterize critical states.
By accurately identifying the invariance of critical states, we can potentially distinguish them from other states in experiments with greater accuracy. Moreover, this insight opens up new possibilities for the precise manipulation of critical states in the future. The ability to manipulate these unique quantum states could have far-reaching implications for various fields, including quantum computing and materials science. This work contributes to advancing our understanding and control over critical states in quantum disordered systems, paving the way for further exploration and potential applications in quantum technologies.
Acknowledgements.
This work was supported by the Natural Science Foundation of Nanjing University of Posts and Telecommunications (Grants No. NY223109).

### v2

@@
 However, accurately characterizing critical states continues to pose a significant challenge.
-In this study, we argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains.
+In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains.
 More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space.
@@
 As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of complex systems.
-Recently, significant advancements have been achieved in the investigation of critical states of wave functions, specifically focusing on critical states of quantum disordered systems Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; gg4 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 .
+Recently, significant advancements have been achieved in the investigation of critical states of quantum disordered systems, specifically focusing on multifractal critical wave functions Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 .
 Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging.
-Nevertheless, the primary challenge in comprehending critical states lies in determining the conditions under which they arise and accurately characterizing them.
+Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and accurately characterizing them.
 Bohr’s complementarity principle MQM provides us with inspiration to bridge this gap.
@@
 The critical state can be considered as a third state in addition to particle and wave properties, thus it can be reasonably inferred that the critical state may exhibit some invariance in both position space and momentum space.
-Figure 1: (Color online) (a) and (b) demonstrate an application of Bohr’s complementarity principle, showing that a localized state in position space must be delocalized in momentum space, and vice versa for an extended state.
+Figure 1: (Color online) Three typical wave functions of disordered systems in dual spaces.
+(a) and (b) demonstrate that a localized state in position space must be delocalized in momentum space, and vice versa for an extended state.
 (c) and (d) illustrate that critical states can simultaneously exhibit delocalization, multifractality, and self-similarity in both position space and momentum space.
@@
 These findings clearly illustrate that critical states maintain a certain kind of invariance across two dual spaces.
+Starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number.
+Contrastively, so far we have not obtained a closed-form solution for the critical state, not even an approximate expression.
+Based on the delocalization, multifractality and self-similarity of critical states, we propose a hypothetical solution, that is, the critical state is a power-law modulated Weierstrass-like function,
+$\psi_{\text{critical}}\sim\sum_{k=0}^{\infty}k^{c}a^{k}\cos(b^{k}\pi n),$ (1)
+where $10<a<$ 1, $b$ is a positive number, and the index $c<0$ is a constant. Thus, we provide a probable expression for the critical state.
 A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces?
-This study demonstrates that the invariance of the critical state is universal, encompassing various physical quantities such as the inverse participation ratio and information entropy.
+This study demonstrates that the invariance of the critical state is universal, encompassing commonly utilized observables in experiments such as the inverse participation ratio and information entropy.
 II Analytical derivation
@@
 This proposed solution provides a reasonable framework encompassing eigenstates of disordered systems and can be utilized to derive mathematical expressions for the Inverse Participation Ratio and information entropy with respect to $\gamma$.
-The Inverse Participation Ratio (IPR) is a physical quantity IPR ; IPR1 ; IPR2 commonly utilized in the fields of quantum mechanics and condensed matter physics to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 , particularly within disordered systems.
-It serves as a quantitative measure of the extent to which a wave function is distributed across space versus its degree of localization.
+The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 .
 A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
-Information entropy $S$, commonly known as Shannon entropy Shannon in honor of Claude Shannon, is a metric for the unpredictability associated with a system.
-Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum complex systems IE1 .
-Information entropy has been utilized to comprehend and quantify the level of unpredictability and randomness in disordered systems IE2 ; IE3 .
+Another typical physical quantity, information entropy $S$, commonly known as Shannon entropy Shannon , is a metric for the unpredictability associated with a system.
+Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 ; IE1 .
+Utilized to quantify the level of unpredictability and uncertainty IE2 ; IE3 , a large $S$ indicates a high unpredictability, corresponding to extended states; while a low $S$ reflects a high predictability, corresponding to localized states.
 The physical quantities IPR and $S$ can be explicitly expressed using the wave function Ansatz $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$.
@@
 Fortunately, diagonalizing the Hamiltonian matrix of the QNE model in momentum space results in a linear eigenvalue problem, significantly simplifying the numerical simulation process.
-Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of critical states, IPR and $S$ in position space and momentum space for the AAH model, respectively.
-(d), (e) and (f) demonstrate the comparison of critical states, IPR and $S$ in position space and momentum space for the QNE model, respectively.
+Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the AAH model, respectively.
+(d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the QNE model, respectively.
 Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space.
@@
 IV Summary
-In conclusion, we have addressed a long-standing unresolved issue regarding the definition of a critical state and its accurate characterization.
+In conclusion, we propose an Ansatz for the critical state, and address a long-standing unresolved issue regarding the definition of a critical state and its accurate characterization.
 Our findings emphasize the universality of critical state invariance across various domains in physics.

### v3

@@
-Critical states exhibit invariance in both position and momentum spaces
+Exactly characterizing critical states: invariance in both position and momentum spaces
 Abstract
-The critical states of disordered systems are intriguing subjects within the realm of condensed matter physics and complex systems.
-These states manifest in materials where disorder plays a significant role, and are distinguished by their multifractal structure and self-similarity.
-However, accurately characterizing critical states continues to pose a significant challenge.
-In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains.
-More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space.
-Subsequent numerical simulations validate the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
+The critical state of disordered systems is an intriguing and popular topic in condensed matter physics, known for its non-ergodic properties such as multifractal structure and self-similarity.
+However, exactly characterizing critical states continues to pose a significant challenge so far.
+In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit invariance in both position and momentum spaces, leading to their delocalization in both dual spaces.
+Furthermore, keeping pace with the idea of Liu-Xia criterion, we demonstrate that physical quantities being directly observed in experiments, such as the inverse participation ratio and information entropy, also exhibit invariance in both position and momentum spaces.
+Subsequent numerical simulations verify the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
 I Introduction
-The concept of critical states Critical in various physical systems pertains to the behaviors and characteristics exhibited during phase transitions or as they approach critical points.
-The examination of critical states provides profound insights into the nature of order, disorder, and collective behavior across diverse disciplines Sachdev1 .
+Critical states, also known as multifractal states, in various physical systems Critical pertains to the characteristics exhibited by phase transitions or across critical points.
+The examination of critical states provides profound insights into the nature of order, disorder, and collective behavior Sachdev1 .
 These states are fundamental for comprehending a wide range of phenomena in statistical physics Critical0 , condensed matter physics Critical1 , and complex systems Critical2 .
@@
 Critical fluctuations have been observed in various systems, including phase transitions in ferromagnets-antiferromagnets Critical4 , percolation thresholds in complex networks Critical5 , and photonic quasicrystals photonic .
-As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of complex systems.
-Recently, significant advancements have been achieved in the investigation of critical states of quantum disordered systems, specifically focusing on multifractal critical wave functions Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 .
+As research continues to advance, the exploration of critical states holds promise for unveiling new phenomena and deepening our understanding of Anderson localization Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; zhou ; Biddle ; Zhang .
+Recently, significant advancements have been achieved in the investigation of critical states in quantum disordered systems, specifically focusing on multifractal wave functions q1 ; q2 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg9 .
 Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging.
-Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and accurately characterizing them.
+Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and exactly characterizing them.
 Bohr’s complementarity principle MQM provides us with inspiration to bridge this gap.
-According to the complementarity principle, it is impossible to simultaneously observe the wave nature and particle nature of matter under different representations, such as position space and momentum space.
-Matter exhibits different properties in these representations.
-The relationship between position space and momentum space is primarily established through the Fourier transform; position space serves as a mathematical framework for describing the location of particles or objects in physical space, while momentum space is perceived as the duality of position space QM .
+According to the complementarity principle, it is impossible to simultaneously observe the wave nature and particle nature of matter under different representations, such as position (or real) space and momentum (or dual) space.
+The relation between position space and momentum space is primarily established through the Fourier transform; position space serves as a mathematical framework for describing the location of particles or objects in physical space, while momentum space is perceived as the duality of position space QM .
+Matter generally exhibits different localization properties in two mutually dual spaces.
 The critical state can be considered as a third state in addition to particle and wave properties, thus it can be reasonably inferred that the critical state may exhibit some invariance in both position space and momentum space.
@@
 (c) and (d) illustrate that critical states can simultaneously exhibit delocalization, multifractality, and self-similarity in both position space and momentum space.
-Based on this duality perception, Liu and Xia propose an explicit criterion Liu-Critical for precisely characterizing critical states.
-They assert that the Lyapunov exponents of critical states should simultaneously be 0 in both position space and momentum space.
-The criterion definitively distinguishes critical states from extended and localized states, as the Lyapunov exponents of the latter two are not equal in position space and momentum space.
+Based on this duality perception, Liu and Xia propose an explicit criterion Liu-Critical for precisely characterizing critical states, asserting that the Lyapunov exponents of critical states should simultaneously be 0 in both position and momentum space.
+The criterion definitively distinguishes critical states from extended and localized states, as the Lyapunov exponents of the latter two are not equal in position and momentum space.
 If one exponent is 0, then the dual exponent must be greater than 0.
-In physics, this also implies that extended and localized states cannot simultaneously exhibit delocalization in position space and momentum space, as illustrated in Fig.
+In more physical language, extended and localized states cannot simultaneously exhibit delocalization in position and momentum space, as illustrated in Fig.
 1(a) and (b).
-In contrast, critical states should demonstrate delocalization, multifractality, and self-similarity in both position space and momentum space, as depicted in Fig.
+In contrast, Liu-Xia criterion predicts that critical states should exhibit delocalization, multifractality, and self-similarity in both position and momentum space, as depicted in Fig.
 1(c) and (d).
+Hence, starting from a strict mathematical formalization perspective, the Lyapunov exponent $\gamma$ in position space and $\gamma_{m}$ in momentum space can be utilized to exactly distinguish extended, localized and critical eigenstates in quantum disordered systems, which are summarized in Table 1.
 These findings clearly illustrate that critical states maintain a certain kind of invariance across two dual spaces.
-Starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number.
+Table 1: Classification of three typical eigenstates in quantum disordered systems according to Liu-Xia criterion.
+Eigenstates Position Momentum Delocalization
+Extended $\gamma=0$ $\gamma_{m}>0$ Position
+Localized $\gamma>0$ $\gamma_{m}=0$ Momentum
+Critical $\gamma=0$ $\gamma_{m}=0$ Both
+Beyond those, starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number.
 Contrastively, so far we have not obtained a closed-form solution for the critical state, not even an approximate expression.
@@
 where $10<a<$ 1, $b$ is a positive number, and the index $c<0$ is a constant. Thus, we provide a probable expression for the critical state.
-A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces?
-This study demonstrates that the invariance of the critical state is universal, encompassing commonly utilized observables in experiments such as the inverse participation ratio and information entropy.
+Unfortunately, the critical state has not been fully confirmed by experiments so far.
+A question arises: can we utilize Liu-Xia criterion to convincingly observe critical states in experiments?
+One shortcoming is that it is not possible to directly measure the Lyapunov exponent in experiments.
+Therefore, the properties of physical quantities that can be directly measured in experiments such as the inverse participation ratio and information entropy in position space and momentum space are the focus of this work.
+Our study demonstrates that the invariance of the critical state is universal, the physical quantities that can be directly observed in these experiments are also invariant in both position and momentum space.
 II Analytical derivation
@@
 For a given disordered potential, the conjectured solution for the wave function in the $n$-th lattice is $|\psi_{n}|\equiv|\psi_{0}|e^{-\gamma n}$, where $\gamma\geq 0$ represents the Lyapunov exponent.
-A value of $\gamma=0$ indicates that $|\psi_{n}|\sim|\psi_{0}$, corresponding to an extended state; while a value of $\gamma>0$ indicates that $|\psi_{n}|\sim|\psi_{0}|e^{-\gamma n}$, corresponding to a localized state.
+A value of $\gamma=0$ indicates that $|\psi_{n}|\sim|\psi_{0}|$, corresponding to an extended state; while a value of $\gamma>0$ indicates that $|\psi_{n}|\sim|\psi_{0}|e^{-\gamma n}$, corresponding to a localized state.
 This proposed solution provides a reasonable framework encompassing eigenstates of disordered systems and can be utilized to derive mathematical expressions for the Inverse Participation Ratio and information entropy with respect to $\gamma$.
-The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 .
+The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions gg1 ; gg2 .
 A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
 Another typical physical quantity, information entropy $S$, commonly known as Shannon entropy Shannon , is a metric for the unpredictability associated with a system.
-Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 ; IE1 .
+Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 .
 Utilized to quantify the level of unpredictability and uncertainty IE2 ; IE3 , a large $S$ indicates a high unpredictability, corresponding to extended states; while a low $S$ reflects a high predictability, corresponding to localized states.
@@
 $S=-\lim_{n\rightarrow\infty}\sum_{n=1}^{L}|\psi_{n}|^{2}\ln(|\psi_{n}|^{2})\sim\gamma+\gamma\coth(\gamma)-\ln(-1+e^{2\gamma}).$ (3)
-According to the Liu-Xia criterion Liu-Critical , critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space.
-This can be mathematically expressed as the equality of the Lyapunov exponent of position space $\gamma$ and that of momentum space $\gamma_{m}$,
+According to the Liu-Xia criterion Liu-Critical , we have
 $\gamma=\gamma_{m}.$ (4)
@@
 ${\rm IPR}={\rm IPR_{m}},~{}~{}~{}~{}S=S_{m}.$ (5)
-Equation (5) represents the key innovation of this study, demonstrating the universal invariance of the critical state in both position space and momentum space.
+Equation (5) demonstrates the universal invariance of the critical state in both position space and momentum space.
 This concept applies not only to Lyapunov exponent, but also to other fundamental physical quantities such as IPR and $S$.
@@
 $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity).
-Thus, the potential $V_{n}$ at each lattice site $n$ is quasiperiodic because it is a non-repeating function with long-range correlations, unlike random disordered potentials that are purely random at each site.
+Thus, the potential $V_{n}$ is quasiperiodic because it is a non-repeating function with long-range correlations, unlike random disordered potentials that are purely random at each site.
 The AAH model yields an intriguing result in the form of a phase transition between two distinct phases as the strength of the quasiperiodic potential $V$ is varied.
@@
 (6)], particularly on the right-hand side of the equation, making it no longer a linear eigenvalue problem like previous quasiperiodic systems.
-The QNE model holds a unique position in the field of Non-Hermitian physics NH1 ; NH2 ; NH3 ; NH4 ; NH5 ; NH6 ; NH7 , because non Hermitian can be introduced in three ways: through non-reciprocity in hopping terms, through complex momentum involving gain and loss of reciprocal hopping terms, or by introducing complex mass representing gain and loss of on-site potential.
+The QNE model holds a unique position in the field of Non-Hermitian physics NH1 ; NH3 ; NH4 ; NH5 ; NH6 ; NH7 , because non Hermitian can be introduced in three ways: through non-reciprocity in hopping terms, through complex momentum involving gain and loss of reciprocal hopping terms, or by introducing complex mass representing gain and loss of on-site potential.
 However, the QNE model does not fall into any of these three cases as it is not a linear eigenvalue problem.
@@
 Fortunately, diagonalizing the Hamiltonian matrix of the QNE model in momentum space results in a linear eigenvalue problem, significantly simplifying the numerical simulation process.
-Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the AAH model, respectively.
-(d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the QNE model, respectively.
+Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the middle eigenstate of AAH model, respectively.
+(d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the $E=0.4$ eigenstate of QNE model, respectively.
 Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space.
@@
 We perform direct numerical diagonalization on the AAH model and QNE model, obtaining the eigenvalues and eigenstates separately.
-The total number of lattices in the system is set to 1000.
-For the AAH model, our focus is on the middle eigenstate; whereas for the QNE model, we concentrate on the eigenstate with an eigenvalue of 0.
+The total number of lattices in the system is set to 987.
+For the AAH model, our focus is on the middle eigenstate; whereas for the QNE model, we concentrate on the eigenstate with an eigenvalue of $E=0.4$.
 The numerical results are summarized in Fig.